OpenStudy (anonymous):

What is the solution for x in the proportion 14 over 18 equals x over 63.? Describe your steps.

OpenStudy (agent0smith):

\[\Large \frac{ 14 }{18 } = \frac{ x }{63 }\] Do you know how to cross-multiply over the equals sign?

OpenStudy (sanjanap):

14/18= x/63 (14/18)(63/1)=x 49

OpenStudy (agent0smith):

Or, just multiply both sides by 63.

OpenStudy (sanjanap):

BAM!!

OpenStudy (sanjanap):

just like @agent0smith said, cross multiply...

OpenStudy (anonymous):

haha, thank you! :)

OpenStudy (agent0smith):

You understand? BAM!!

OpenStudy (anonymous):

yes :)

OpenStudy (sanjanap):

HAHAHA @agent0smith

OpenStudy (agent0smith):

:D